Ben is 12 years older than Ishaan. Ben and Ishaan first met two years ago. Three years ago, Ben was 4 times as old as Ishaan.
How old is Ishaan now?

Found 2 solutions by josgarithmetic, Edwin McCravy:
Answer by josgarithmetic(39616)
Answer by Edwin McCravy(20054)


Ben is 12 years older than Ishaan.
B = I + 12

Ben and Ishaan first met two years ago.
That is irrelevant information, i.e., that has nothing whatever 
to do with this problem.

Three years ago, Ben was 4 times as old as Ishaan.
B-3 = 4(I-3)
B-3 = 4I-12
  B = 4I-9

4I - 9 = I + 12
    3I = 21
     I = 7

How old is Ishaan now?

7

Edwin